OpenStudy (anonymous):

The measure of angle 6 is 32. What is m<4+m<5?

OpenStudy (anonymous):



OpenStudy (anonymous):

Angle 6 is supplementary to Angle 5, so they add to 180. Also, Angle 5 and Angle 4 are equal, since they are exterior alternate angles. so m<6 + m<5 = 180 32 + m<5 = 180 m<5 = 180 - 32 = 148 and m<4 = m<5 = 148 So m<4 + m<5 = 148 + 148